# Symplectic mapping class relations from pencil pairs

### v1

@@ 2.3. Some contact-topological motivations
We outline the role of positive factorizations and positive relations in modern contact topology. This is an advertisement aiming at those readers who already have some exposure to contact topology and so the exposition will be light on details.
Given a positive factorization $\vec{L}$ as above we can build a Liouville manifold $W=W_{\vec{L}}$ of $\dim W_{\vec{L}}=\dim F+2$ as follows: Take $W_{0}$ to be $(\mathbb{D}\times F,r^{2}d\theta+\beta_{F})$ with the corner $\partial\mathbb{D}\times\partial F$ rounded. The $\Lambda_{j}=(e^{j2\pi i/m},L_{j})$ will be framed Legendrian spheres in the contact boundary of $W_{0}$ (after a slight perturbation of the boundary of $W_{0}$ to manage the exact $1$-forms $\beta_{F}|_{TL_{i}}$). Then $W$ is obtained by attaching $\operatorname{ind}=\frac{1}{2}\dim W_{0}$ Weinstein handles along the $\Lambda_{j}$. It follows that
@@
When $(F,\beta_{F})$ is Weinstein, it follows that $(W,\beta_{W})$ is as well. In any case, the boundary of $W$ is naturally a contact manifold $(\Gamma,\xi_{\Gamma})$ described as an open book $\mathcal{OB}(F,\beta_{F},\phi)$ with page $(F,\beta_{F})$ and monodromy $\phi$, whose contact structure depends only on $\phi\in\pi_{0}\operatorname{Symp}^{c}(F,d\beta_{F})$ [BHH23, Gi02, TW75]. It’s known that any Weinstein manifold of $\dim\geq 4$ can be constructed in this way with $(F,\beta_{F})$ Weinstein [BHH23, GP17].
It follows that a positive relation $\vec{L},\vec{L}^{\prime}$ gives us two possibly distinct Liouville fillings $(W=W_{\vec{L}},\beta)$ and $(W^{\prime}=W_{\vec{L}^{\prime}},\beta^{\prime})$ of the same $(\Gamma,\xi_{\Gamma^{\prime}})$. So positive relations help us to construct distinct Liouville and Weinstein fillings of a given contact manifold, cf. [O18, OS04a].
Provided a positive relation $\vec{L},\vec{L}^{\prime}$, the presentation of the contact boundaries of $W$ and $W^{\prime}$ as open books with the same monodromy give us an identifications of these boundaries. As described in [Br23], we can use these fillings to determine a $\tau$-invariant contact structure $\xi$ on $N=(-\epsilon,\epsilon)_{\tau}\times S,S=W\cap_{\Gamma}W^{\prime}$. We say that $S$ is a convex hypersurface with standard neighborhood $(N,\xi)$. Identifying $S=\{0\}\times S$, the Euler class of $\xi|_{S}$ is computed
@@
Combining [HH19] and [Br23], any oriented hypersurface in a contact manifold $(M,\xi)$ of $\dim M\geq 5$ has such a neighborhood (determined by a positive relation on a Weinstein $(F,\beta_{F})$) after applying a $\mathcal{C}^{0}$ small perturbation to the hypersurface. Since convex hypersurfaces exist in abundance by [HH19], so should positive relations. Moreover, Equation (2.3.1) tells us that there are lots of positive relations $\vec{L},\vec{L}^{\prime}$ which are not braid like. Here are some examples.
@@
For a fixed $n\in\mathbb{Z}_{>2}$ and a partition $\vec{n}=(n_{1},\cdots,n_{m}),n_{i}\leq n_{i+1},\sum n_{i}=n$, let $\mathbb{P}^{\vec{n}}=\mathbb{P}^{n_{1}}\times\cdots\times\mathbb{P}^{n_{m}}$. Then $T\mathbb{P}^{\vec{n}}$ defines an almost contact structure on the $\dim=2n+1$ manifold $M^{\vec{n}}=\mathbb{S}^{1}\times\mathbb{P}^{\vec{n}}$. By [BEM15] – which includes a definition of almost contact structures – there is a (genuine) contact structure $\xi^{\vec{n}}$ on $M^{\vec{n}}$ which is homotopic to $T\mathbb{P}^{\vec{n}}\subset TM^{\vec{n}}$ through almost contact structures.
Let $(M,\xi)$ be the contact connected sum of the $(M^{\vec{n}},\xi^{\vec{n}})$, over all partitions $\vec{n}$, the number of which we’ll write as $N_{n}$. For $\vec{r}=(r_{\vec{n}})\in\mathbb{Z}^{N_{n}}$ consider the homology class $h_{\vec{r}}\in H_{2n}(M)$ given by the sum of the $r_{\vec{n}}[\{\theta\}\times\mathbb{P}^{\vec{n}}]$, with each summand living in the image of $H_{2n}(M^{\vec{n}})\simeq H_{2n}(M^{\vec{n}}\setminus\mathbb{D}^{2n+1})\rightarrow H_{2n}(M)$. It’s not difficult to see that $h_{\vec{r}}$ has a connected, embedded, oriented representative, $S_{0}$. By [HH19] we can apply a $\mathcal{C}^{0}$-small perturbation to $S_{0}$ to obtain a $S\subset M$ which is convex with neighborhood $(N,\xi)$. The contact structure $\xi|_{S}$ gives $S$ the structure of a stably almost complex manifold whose cobordism class is $[S]=\sum r_{\vec{n}}\mathbb{P}^{\vec{n}}\in\Omega^{U}_{2n}$. Since the $\mathbb{P}^{\vec{n}}$ additively generate $\Omega^{U}_{2n}$ (cf. [MiS74]), we see that any almost complex cobordism class can be represented by a convex hypersurface.
By [Br23] there is some Weinstein $(F,\beta_{F})$ of $\dim=2n-2$ having a positive relation $\vec{L},\vec{L}^{\prime}$ determining $(N,\xi)$. Computing $(-1)^{n}\int_{S}e(\xi)=(-1)^{n}\int_{S}c_{n}(\xi)$ gives us the difference in the lengths of the positive relation. Note that the image of $\int_{\ast}c_{n}:\Omega^{U}_{2n}\rightarrow\mathbb{Z}$ is $2\mathbb{Z}$ when $n$ is odd and $\mathbb{Z}$ when $n$ is even. Taking $m=n-1$, we see that for every $m\geq 2$ there are Weinstein $(F,\beta_{F})$ of $\dim=2m$ having positive relations with $|\vec{L}|-|\vec{L}^{\prime}|$ being any integer when $m$ is odd and any even integer when $m$ is even.
Unfortunately it is very difficult to explicitly describe such $(N,\xi)$, even for convex hypersurfaces in well-studied $(M,\xi)$. In searching for symplectic mapping class relations, we seek to have some interesting examples of convex hypersurfaces at least for which the $(F,\beta_{F})$ can be explicitly described (eg. as affine varieties). In a follow-up article, we will study the $(N,\xi)$ determined by some of the positive relations in this article using the techniques of [Av23a].
@@ 2.4. Review of algebraic Lefschetz pencils
Let $\operatorname{\mathcal{L}}$ be an ample line bundle on a closed variety $X$. The pair $(X,\operatorname{\mathcal{L}})$ – a polarization of $X$ – determines an immersion $\phi_{\operatorname{\mathcal{L}}}:X\rightarrow\mathbb{P}^{m}$, $m=\dim_{\mathbb{C}}H^{0}(X,\operatorname{\mathcal{L}})-1$ such that $\phi_{\operatorname{\mathcal{L}}}^{\ast}\mathcal{O}_{1}=\operatorname{\mathcal{L}}$ and for which $\omega_{\operatorname{\mathcal{L}}}=\phi_{\operatorname{\mathcal{L}}}^{\ast}\omega_{\mathbb{P}^{m}}$ is symplectic. Here $\omega_{\mathbb{P}^{m}}$ is the Fubini-Study form, normalized to have volume $1$. When $\phi_{\operatorname{\mathcal{L}}}$ is an embedding, $\operatorname{\mathcal{L}}$ is very ample. Replacing $\operatorname{\mathcal{L}}$ with $\operatorname{\mathcal{L}}^{\otimes k}$ for $k\in\mathbb{Z}_{>0}$ gives more ample line bundles.
@@
    subvarieties $Z_{i}=\{\boldsymbol{\mathfrak{s}}_{i}=0\}\subset X$ (ie. an intersection of $X\subset\mathbb{P}^{m}$ with a hyperplane determined by a section of $\mathcal{O}_{1}\rightarrow\mathbb{P}^{m}$) with $\operatorname{expdim}_{\mathbb{C}}V_{i}=\dim_{\mathbb{C}}X-1$,
@@
    a base locus $B=B_{\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1}}=Z_{0}\cap Z_{1}$ with $\operatorname{expdim}_{\mathbb{C}}B=\dim_{\mathbb{C}}X-2$, and
@@
    The $Z_{i}$ are transversely cut out, so that $\operatorname{PD}_{X}(\omega_{\operatorname{\mathcal{L}}})=[Z_{i}]\in H_{\dim X-2}(X)$,
@@
It follows from Bertini’s theorem that a generic pair of sections of an ample bundle will determine a pencil. The proof of the following lemma is well-known. We give a quick sketch since variations of the argument will be used throughout the article.
@@
Consider a generic $T\in[0,1]$ family $(\boldsymbol{\mathfrak{s}}_{0}^{T},\boldsymbol{\mathfrak{s}}_{1}^{T})$ of pairs of sections of $\operatorname{\mathcal{L}}$. By Bertini’s theorem the set of such pairs for which either $\{\boldsymbol{\mathfrak{s}}_{0}=0\}$ or $\{\boldsymbol{\mathfrak{s}}_{0}=\boldsymbol{\mathfrak{s}}_{1}=0\}$ has complex codimension $\geq 1$ (real codimension $\geq 2$), so we can assume that the path of pairs of sections determines a path of pencils. Applying a Darboux-Moser-Weinstein argument [MS99] to the $T$-family establishes the claim. ∎
When $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a Lefschetz pencil then $\pi$ is a Lefschetz fibration. We will outline the essential properties and point to the references, eg. [AS08, D96b, OS04b, Se08], for a full definition. The complement
@@
is an affine variety and so is naturally a Weinstein manifold whose underlying symplectic form is $\omega_{\operatorname{\mathcal{L}}}$. So are the regular fibers $F_{z}:=\pi^{-1}(z),z\in\mathbb{C}\setminus\operatorname{Critv}(\pi)$ of $\pi$. The closure of each $F_{z}$ in $X$ intersects $Z_{0}$ along $B$. By considering holonomy along paths in $\mathbb{C}\setminus\operatorname{Critv}(\pi)$, each $F_{z}$ is deformation equivalent to
$F:=F_{0}=Z_{1}\setminus B.$
The pencil $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ determines a compactly supported symplectomorphism of $(F,\omega_{\operatorname{\mathcal{L}}})$ in two ways. First, we study the monodromy of $\pi$. Let $\eta_{i}$ be a collection of embedded, oriented, $[0,1]$-parameterized paths in $\mathbb{C}$ indexed by the $z_{i}\in\operatorname{Critv}(\pi)$, begin at $0\in\mathbb{C}$ and ending at $z_{i}$, and disjoint away from $0\in\mathbb{C}$. We require that the $i$ are ordered so that tangent vectors of the $\eta_{i}$ at 00 are ordered counterclockwise. Each $\eta_{i}$ determines a Lagrangian disk $D_{i}$ in $W$ for which $\pi D_{i}=\eta$ and for which $L_{i}=\partial D_{i}$ is an exact Lagrangian sphere in $F$. The $\eta_{i}$ are vanishing paths and the $D_{i}$ are their associated Lefschetz thimbles which provide framings of the $L_{i}$. Let $\gamma\subset\mathbb{C}\setminus\operatorname{Critv}(\pi)$ be a parameterized simple closed curve based at 00 and oriented counterclockwise, otherwise disjoint from the $\eta_{i}$, encircling all of the $\eta_{i}$, and encircling $\eta_{1}$ first. Then the monodromy morphism $\Phi_{\gamma}\in\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$ can be expressed as a product of Dehn twists $\Phi_{\gamma}=\tau_{L_{\#\operatorname{Crit}(\pi)}}\cdots\tau_{L_{1}}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega)$, generalizing [Arn95]. Modifying the choices of $(\eta_{i})$ corresponds to application of conjugacy relations on the $\vec{L}$.
Second, note that a collar neighborhood of the ideal boundary $\Gamma$ of $F$ is identical to a $\mathbb{D}$-bundle neighborhood of $B\subset Z_{1}$ with its zero section removed. So we can associate to this data a fibered Dehn twist $\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$, cf. [CDvK14, O21]. We give a brief review: Our collar neighborhood of $F$ is given as a symplectization $([0,\infty)_{s}\times\Gamma,e^{s}\alpha)$ determined by a contact form $\alpha$ on $\Gamma$ whose Reeb flow $R_{\alpha}$ is $2\pi$-periodic. Using this collar we can write
@@
for a smooth function $g=g(s)$ which is increasing with $g(0)=0$ and $g(s)=2\pi$ for $s\geq 1$. Then $\tau_{\partial F}$ extends to the rest of $F$ as the identity map. The induced action on $H_{\ast}(F)$ by $\tau_{\partial F}$ is then clearly the identity. Observe that when $\dim F=2$, $\tau_{\partial F}$ always exists and is a product of positive Dehn twists along simple curves, one parallel to each boundary component of $F$. These simple closed curves are smoothly isotopic to exact Lagrangians iff $F$ has at least two boundary components.
The first statement of following theorem is proved in [Aro01, Go04], with special cases worked out in [AA16, O21]. The count of critical points as a weighted sum of Euler characteristics follows easily from the Lefschetz hyperplane theorem, cf. [O21, Theorem 4.5].
@@
In particular, $\tau_{\partial F_{0}}$ is isotopic to a symplectomorphism determined by a product of
(2.4.1) $|\vec{L}|=\#\operatorname{Crit}\pi=(-1)^{\dim_{\mathbb{C}}X}\left(\chi(X)-2\chi(Z_{1})+\chi(B)\right)$
@@
It’s easy to see that not all $\tau_{\partial F}$ can be factorized as products of Dehn twists along exact Lagrangian spheres. For example, take $(F,\beta_{F})$ to be the complement of a $\deg\geq 2$ hypersurface in $\mathbb{P}^{n},n>1$. The ideal boundary of $F$ is a $\mathbb{S}^{1}$-fibered contact manifold so that $\tau_{\partial F}$ is defined. However $F$ contains no Lagrangian spheres, since $\mathbb{P}^{n}$ contains no Lagrangian spheres – a fact which is obvious for homological reasons when $n$ is even and non-trivial when $n$ is odd [Se00]. Addition examples due to Biran and Giroux are described in [CDvK14].

### v2

@@ 2.3. Some contact-topological motivations
We outline the role of positive factorizations and positive relations in modern contact topology. As stated in the introduction, our goal is to demonstrate how recent advances in the field [Br23, BHH23, HH19] imply the existence of many positive relations on high-dimensional Weinstein manifolds. This subsection is aimed at those readers who already have some exposure to contact topology, so our exposition will be light on details.
Given a positive factorization $\vec{L}$ as above we can build a Liouville manifold $W=W_{\vec{L}}$ of $\dim W_{\vec{L}}=\dim F+2$ as follows: Take $W_{0}$ to be $(\mathbb{D}\times F,r^{2}d\theta+\beta_{F})$ with the corner $\partial\mathbb{D}\times\partial F$ rounded. The $\Lambda_{j}=(e^{j2\pi i/m},L_{j})$ will be framed Legendrian spheres in the contact boundary of $W_{0}$ (after a slight perturbation of the boundary of $W_{0}$ to manage the exact $1$-forms $\beta_{F}|_{TL_{i}}$). Then $W$ is obtained by attaching $\operatorname{ind}=\frac{1}{2}\dim W_{0}$ Weinstein handles along the $\Lambda_{j}$. Such a $(W,\beta)$ is naturally the total space of a Lefschetz fibration, as will be discussed in the next subsection. It follows that
@@
When $(F,\beta_{F})$ is Weinstein, $(W,\beta_{W})$ is as well. In any case, the boundary of $W$ is naturally a contact manifold $(\Gamma,\xi_{\Gamma})$ described as an open book with page $(F,\beta_{F})$ and monodromy $\phi$ a product of the $\tau_{L_{i}}$. This contact structure depends only on $\phi\in\pi_{0}\operatorname{Symp}^{c}(F,d\beta_{F})$ [BHH23, Gi02, TW75]. According to [BHH23, GP17] any Weinstein manifold of $\dim\geq 4$ can be constructed in this way with $(F,\beta_{F})$ Weinstein.
It follows that each positive relation $\vec{L},\vec{L}^{\prime}$ gives us two possibly distinct Liouville fillings $(W=W_{\vec{L}},\beta)$ and $(W^{\prime}=W_{\vec{L}^{\prime}},\beta^{\prime})$ of the same contact manifold $(\Gamma,\xi_{\Gamma^{\prime}})$. Consequently, positive relations help us to construct distinct Liouville and Weinstein fillings of a given contact manifold, cf. [O18, OS04a]. For an alternate construction of distinct fillings using relations in symplectomorphism groups, see [Z24].
Provided a positive relation $\vec{L},\vec{L}^{\prime}$, the presentation of the contact boundaries of the Liouville manifolds $W$ and $W^{\prime}$ as open books with the same monodromy give us an identifications of these boundaries. As described in [Br23], we can use the fillings $W,W^{\prime}$ to determine a $(-\epsilon,\epsilon)$-invariant contact structure $\xi$ on
$N=(-\epsilon,\epsilon)\times S,\quad S=W\cap_{\Gamma}W^{\prime}.$
We say that $S$ is a convex hypersurface with standard neighborhood $(N,\xi)$. Identifying $S=\{0\}\times S$, the Euler number of $\xi|_{S}$ is computed
@@
The example below demonstrates that the only constraints on the existence of a $(N,\xi)$ determining a fixed Euler number is unconstrained, beyond a cobordism-theoretic restriction that it must be even for $n$ even.
By [HH19, Theorem 1.2.3], any oriented hypersurface in a contact manifold $(M,\xi)$ of $\dim M\geq 5$ can be $\mathcal{C}^{0}$-approximated by a convex hypersurface $S\subset M$ whose $W$ and $W^{\prime}$ are Weinstein. Moreover [Br23, Theorem 1.8] states that such an $S$ can be assumed to be the result of the above construction determined by a positive relation. Combining [HH19] and [Br23] we see that convex hypersurfaces exist in abundance, as do positive relations. Moreover, Equation (2.3.1) tells us that there are lots of positive relations $\vec{L},\vec{L}^{\prime}$ which are not braid-like. Here are some examples.
@@
For a fixed $n\in\mathbb{Z}_{\geq 2}$ and a partition $\vec{n}=(n_{1},\dots,n_{m})$, $n_{i}\leq n_{i+1}$, $\sum n_{i}=n$, write
$\mathbb{CP}^{\vec{n}}=\mathbb{CP}^{n_{1}}\times\cdots\times\mathbb{CP}^{n_{m}},\quad M^{\vec{n}}=\mathbb{S}^{1}\times\mathbb{CP}^{\vec{n}}$
The complex structure on $T\mathbb{CP}^{\vec{n}}$ defines an almost contact structure on the $\dim=2n+1$ manifold $M^{\vec{n}}$. By [BEM15] – which includes a definition of almost contact structures – there is a (genuine) contact structure $\xi^{\vec{n}}$ on $M^{\vec{n}}$ which is homotopic to $T\mathbb{CP}^{\vec{n}}\subset TM^{\vec{n}}$ through almost contact structures.
Let $(M,\xi)$ be the contact connected sum of the $(M^{\vec{n}},\xi^{\vec{n}})$, over all partitions $\vec{n}$, the number of which we’ll write as $N_{n}$. For $\vec{r}=(r_{\vec{n}})\in\mathbb{Z}^{N_{n}}$ consider the homology class $h_{\vec{r}}\in H_{2n}(M)$ given by the sum of the $r_{\vec{n}}[\{\theta\}\times\mathbb{CP}^{\vec{n}}]$, with each summand living in the image of the composition
$H_{2n}(M^{\vec{n}})\xrightarrow{\simeq}H_{2n}(M^{\vec{n}}\setminus\mathbb{D}^{2n+1})\rightarrow H_{2n}(M).$
It’s not difficult to see that $h_{\vec{r}}$ has a connected, embedded, oriented representative, $S_{0}$. By [HH19] we can apply a $\mathcal{C}^{0}$-small perturbation to $S_{0}$ to obtain an orientable, $\operatorname{codim}=1$ submanifold $S\subset M$ which is convex with neighborhood $(N,\xi)$. The contact structure $\xi|_{S}$ gives $S$ the structure of a stably almost complex manifold whose cobordism class is $[S]=\sum r_{\vec{n}}\mathbb{CP}^{\vec{n}}\in\Omega^{U}_{2n}$. Since the $\mathbb{CP}^{\vec{n}}$ additively generate $\Omega^{U}_{2n}$ (cf. [MiS74]), we see that any almost complex cobordism class can be represented by a convex hypersurface.
By [Br23, Theorem 1.8] there is some Weinstein $(F,\beta_{F})$ of $\dim=2n-2$ having a positive relation $\vec{L},\vec{L}^{\prime}$ determining $(N,\xi)$. By Equation (2.3.1), $\int_{S}e(\xi)=\int_{S}c_{n}(\xi)$ determines the difference in the lengths of the positive factorizations. Note that the image of $\int_{\ast}c_{n}:\Omega^{U}_{2n}\rightarrow\mathbb{Z}$ is $2\mathbb{Z}$ when $n$ is odd and $\mathbb{Z}$ when $n$ is even. Taking $m=n-1$, we see that for every $m\geq 2$ there are Weinstein $(F,\beta_{F})$ of $\dim=2m$ having positive relations with $|\vec{L}|-|\vec{L}^{\prime}|$ being any integer when $m$ is odd and any even integer when $m$ is even.
As convex hypersurfaces $S$ are only $\mathcal{C}^{0}$-dense, rather than $\mathcal{C}^{1}$-dense in $\dim\geq 5$ contact manifolds (cf. [Ch24]), it is very difficult to explicitly describe such convex surface neighborhoods $(N\supset S,\xi)$, even for convex hypersurfaces in well-studied $(M,\xi)$. In searching for symplectic mapping class relations, we seek to have some interesting examples of convex hypersurfaces at least for which the $(F,\beta_{F})$ can be described somewhat explicitly (e.g. as affine or quasi-projective varieties). In a follow-up article, we will study the $(N,\xi)$ determined by some of the positive relations in this article using the techniques of [Av23a].
@@ 2.4. Review of algebraic Lefschetz pencils
Let $\operatorname{\mathcal{L}}$ be a very ample line bundle on a smooth, closed variety $X$. By definition, this means that a $\mathbb{C}$-basis $(\boldsymbol{\mathfrak{s}}_{i})_{i=0}^{N}$ of $H^{0}(X,\operatorname{\mathcal{L}})$, the space of holomorphic sections of $\operatorname{\mathcal{L}}\to X$, determines a holomorphic embedding
(2.4.1) $\phi_{\operatorname{\mathcal{L}}}:X\rightarrow\mathbb{CP}^{N},\quad\phi_{\operatorname{\mathcal{L}}}(z)=[\boldsymbol{\mathfrak{s}}_{0}(z):\cdots:\boldsymbol{\mathfrak{s}}_{N}(z)],\quad N=\dim_{\mathbb{C}}H_{0}(X,\operatorname{\mathcal{L}})-1.$
A change of basis corresponds to composing $\phi_{\operatorname{\mathcal{L}}}$ with a linear automorphism of $\mathbb{CP}^{N}$, which is a biholomorphism. Since the complement of each hyperplane $\{z_{i}=0\}\simeq\mathbb{CP}^{N-1}\subset\mathbb{CP}^{N}$ is affine, we see that $\phi_{\operatorname{\mathcal{L}}}$ restricted to the complement of each
$Z_{i}=\{\boldsymbol{\mathfrak{s}}_{i}=0\}=(\{z_{i}=0\}\cap X)\subset X$
gives an embedding into $\mathbb{C}^{N}$. So each $X\setminus\{\boldsymbol{\mathfrak{s}}_{i}=0\}$ is naturally an affine variety. Replacing $\operatorname{\mathcal{L}}$ with $\operatorname{\mathcal{L}}^{\otimes k}=\phi_{\operatorname{\mathcal{L}}}^{\ast}\mathcal{O}_{k}$ for $k\in\mathbb{Z}_{>0}$ gives more very ample line bundles on $X$. Since $\phi_{\operatorname{\mathcal{L}}}$ is holomorphic,
$\omega_{\operatorname{\mathcal{L}}}=\phi_{\operatorname{\mathcal{L}}}^{\ast}\omega_{\mathbb{CP}^{N}}\in\Omega^{2}(X),\quad\omega_{\mathbb{CP}^{N}}=\frac{i}{2\pi}\partial\overline{\partial}\log(\left\lVert z\right\rVert^{2})\in\Omega^{1}(\mathbb{CP}^{N})$
is a symplectic form on $X$ with respect to which we can view $\phi_{\operatorname{\mathcal{L}}}$ as a symplectic embedding. This $\omega_{\mathbb{CP}^{N}}$ is the Fubini-Study form, normalized to have volume $1$.
@@
    subvarieties $Z_{i}\subset X$ of $\dim_{\mathbb{C}}Z_{i}=\dim_{\mathbb{C}}X-1$ when the sections $\boldsymbol{\mathfrak{s}}_{i}$ are transversely cut out,
@@
    a base locus $B=Z_{0}\cap Z_{1}$ of $\dim_{\mathbb{C}}B=\dim_{\mathbb{C}}X-2$ when $\boldsymbol{\mathfrak{s}}_{0}\oplus\boldsymbol{\mathfrak{s}}_{1}$ is transversely cut out, and
@@
    The $Z_{i}$ are transversely cut out, and hence smooth with $\operatorname{PD}_{X}(\omega_{\operatorname{\mathcal{L}}})=[Z_{i}]\in H_{\dim X-2}(X)$,
@@
We remark that the transversality requirements in this definition are non-standard. It follows from Bertini’s theorem that a generic pair of sections of an ample bundle will determine a pencil which extend to a $\phi_{\operatorname{\mathcal{L}}}$ as above. See, for example, [GH94, §1.1] for a review. The proof of the following lemma is well-known. We give a quick sketch since variations of the proof will be used throughout the article.
@@
Consider a generic $T\in[0,1]$ family $(\boldsymbol{\mathfrak{s}}_{0}^{T},\boldsymbol{\mathfrak{s}}_{1}^{T})$ of pairs of sections of $\operatorname{\mathcal{L}}$. By Bertini’s theorem the set of such pairs for which $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ fails to be a Lefschetz-pencil has complex codimension $\geq 1$, that is, real codimension $\geq 2$. Therefore we can assume that the path of pairs of sections determines a path of pencils. Applying a Darboux-Moser-Weinstein argument [MS99, §3.2] to the $T$-family establishes the claim. ∎
When $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ is a Lefschetz pencil then $\pi$ is a Lefschetz fibration. We will outline the essential properties and point to the references, e.g. [AS08, D96b, OS04b, Se08, T20], for a full definition. The complement
@@
is a smooth affine variety as described above. Therefore $W$ inherits a Liouville structure from
$\beta=-\frac{i}{2\pi}dh\circ J_{0}\in\Omega^{1}(\mathbb{C}^{N}),\quad h=\log(\left\lVert z\right\rVert^{2}+1)$
A computation yields $d\beta=\omega_{\mathbb{CP}^{N}}$ in an affine chart so that $d\beta|_{TW}=\omega_{\operatorname{\mathcal{L}}}$. The plurisubharmonic function $h$ provided, in fact, presents $\mathbb{C}^{N}$ and $W$ as Stein manifolds, determining Weinstein manifold structures [CE12]. The regular fibers
$F_{z}=\pi^{-1}(z),\quad z\in\mathbb{C}\setminus\operatorname{Critv}(\pi)$
of $\pi$ are then also smooth, affine varieties and hence naturally Weinstein. The closure of each $F_{z}$ in the closure $X$ of $W$ intersects $Z_{0}$ along $B$. By considering lifts of paths in $\mathbb{C}\setminus\operatorname{Critv}(\pi)$ via $\pi^{-1}$, we see that each $(F_{z},\beta)$ is deformation equivalent to
$(F,\beta),\quad F:=F_{0}=Z_{1}\setminus B.$
The pencil $(\boldsymbol{\mathfrak{s}}_{0},\boldsymbol{\mathfrak{s}}_{1})$ determines a compactly supported symplectomorphism of $(F,d\beta)$ in two ways. One way to express this symplectomorphism comes from studying the monodromy of $\pi$. Let $\eta_{i}$ be a collection of embedded, oriented, $[0,1]$-parameterized paths in $\mathbb{C}$ indexed by the $z_{i}\in\operatorname{Critv}(\pi)$, beginning at $0\in\mathbb{C}$, ending at $z_{i}$, and disjoint away from $0\in\mathbb{C}$. We require that the indices $i$ are ordered so that tangent vectors of the $\eta_{i}$ at 00 are ordered counterclockwise. Each $\eta_{i}$ determines a Lagrangian disk $D_{i}$ in $W$ for which $\pi D_{i}=\eta$ and for which $L_{i}=\partial D_{i}$ is an exact Lagrangian sphere in $F$. The $\eta_{i}$ are vanishing paths and the $D_{i}$ are their associated Lefschetz thimbles which provide framings of the $L_{i}$. Let $\gamma\subset\mathbb{C}\setminus\operatorname{Critv}(\pi)$ be a parameterized simple closed curve based at 00 and oriented counterclockwise, otherwise disjoint from the $\eta_{i}$, encircling all of the $\eta_{i}$, and encircling $\eta_{1}$ first. Then the monodromy morphism $\Phi_{\gamma}\in\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$ can be expressed as a product of Dehn twists
$\Phi_{\gamma}=\tau_{L_{\#\operatorname{Crit}(\pi)}}\cdots\tau_{L_{1}}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega),$
generalizing [Arn95]. Modifying the choice of paths $(\eta_{i})$ applies conjugacy relations to $\vec{L}=(L_{i})$.
Second, note that a collar neighborhood of the ideal boundary $\Gamma$ of $F$ is identical to a $\mathbb{D}$-bundle neighborhood of $B\subset Z_{1}$ with its zero section removed. We can associate to this data a fibered Dehn twist $\tau_{\partial F}\in\pi_{0}\operatorname{Symp}^{c}(F,\omega_{\operatorname{\mathcal{L}}})$, cf. [CDvK14, O21]. We give a brief review: Our collar neighborhood of $F$ with a neighborhood of $B$ removed is given as a finite symplectization $([0,1]_{s}\times\Gamma,d(e^{s}\alpha))$. Here $\alpha$ is a contact form on $\Gamma$ whose Reeb flow $R_{\alpha}$ is $2\pi$-periodic, generating circular rotation on the disk bundle. Using this collar we can write
@@
for a smooth function $g=g(s)$ which is decreasing with $g(0)=2\pi$ and $g(s)=0$ for $s\geq 1$. Then $\tau_{\partial F}$ extends to the rest of $F$ as the identity map. The induced action on $H_{\ast}(F)$ by $\tau_{\partial F}$ is then clearly the identity.
The first statement of the following theorem is proved in [Aro01, Go04], with special cases worked out in [AA16, O21]. The count of critical points as a weighted sum of Euler characteristics follows easily from the Lefschetz hyperplane theorem, cf. [O21, Theorem 4.5].
@@
In particular, $\tau_{\partial F}$ is isotopic to a symplectomorphism determined by a product of
$|\vec{L}|=\#\operatorname{Crit}\pi=(-1)^{\dim_{\mathbb{C}}X}\left(\chi(X)-2\chi(Z_{1})+\chi(B)\right)$
@@
Note that $\tau_{\partial F}$ can be defined on any Liouville manifold $(F,\beta_{F})$ whose ideal contact boundary is a Boothby-Wang circle bundle over some $(B,\omega)$. See [CDvK14] for further details. In this general setting, not all such $\tau_{\partial F}$ can be factorized as products of Dehn twists along exact Lagrangian spheres. For example, take $(F,\beta_{F})$ to be the complement of a $\deg\geq 2$ hypersurface in $\mathbb{CP}^{n},n>1$. The ideal boundary of $F$ is of Boothby-Wang type, so that $\tau_{\partial F}$ is defined but such $(F,d\beta_{F})$ contain no Lagrangian spheres since $\mathbb{CP}^{n}$ contains no Lagrangian spheres. This is obvious for homological reasons when $n$ is even and non-trivially follows from [Se00] for $n$ is odd. When $\deg=2$, these $(F,d\beta_{F})$ coincides with $(\mathbb{D}^{\ast}\mathbb{RP}^{n},dp\wedge dq)$ [Bi01]. See [Au07, T20] for generalizations to cotangent disk bundles of other symmetric spaces. A general criteria for finding fibered Dehn twists which cannot by factored into products of Dehn twists along Lagrangian spheres is given in [CDvK14] (attributed to Biran and Giroux). For a generalization of the definition of fibered Dehn twist provided above, see [P07, T20].